OpenStudy (anonymous):

(x−5)^9(x−2)^4(x+10)^8 urggg this is really frustrating, someone post solution please

OpenStudy (anonymous):

where is the function increasing

OpenStudy (anonymous):

woah this is a scary one but you can figure this out by using the powers ... i just don't remember off the top of my head how that works lol

OpenStudy (anonymous):

you know the x-intercepts so you just have to figure out how even functions and odd functions work

OpenStudy (amistre64):

you take the first derivative; its just the product rule

OpenStudy (amistre64):

the zeros remain the same for the derivative; 5, 2, and -10

OpenStudy (amistre64):

then use an advantageous number inbetween them to test for sign

OpenStudy (anonymous):

i tested them and they are increasing on every interval. how do i write that in interval notation

OpenStudy (anonymous):

Increase : \[(-\infty,\infty)\]?

OpenStudy (amistre64):

if its increasing everywhere except where it zeros; you want to Union up the intersections

OpenStudy (amistre64):

(-inf,a)U(a,b)U(b,inf)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D%28x%E2%88%925%29%5E9%28x%E2%88%922%29%5E4%28x%2B10%29%5E8 you might wanna rechk your interval

OpenStudy (anonymous):

(-inf,-10)U(2,5)U(5,inf) ???

OpenStudy (amistre64):

that would be it if your pos/neg was accurate

OpenStudy (anonymous):

is it not? can you check pls?

OpenStudy (amistre64):

(x−5)^9(x−2)^4(x+10)^8 9(x−5)^8 (x−2)^4 (x+10)^8+ (x−5)^9 4(x−2)^3 (x+10)^8 + (x−5)^9 (x−2)^4 8(x+10)^7 wolfram would be simpler

OpenStudy (anonymous):

for the first interval (-inf, -10) 9(-11−5)^8 (-11−2)^4 = 1104

OpenStudy (amistre64):

thats interval is + yes but i keep getting something wierd around x=-5

OpenStudy (amistre64):

your inputing that into f(x); f'(x) defines the slope; and whenever the slope is positive we are increasing

OpenStudy (anonymous):

am i using the derivative or the original equation to determine where the slope is increasing?

OpenStudy (amistre64):

the derivative defines slope; slope tells us increase or decrese

OpenStudy (amistre64):

ive zoomed in at 5 points; x=-10,-4.75,2,3,5 these seem to be zeros

OpenStudy (amistre64):